# The Bayesian Geometry of Transformer Attention

### v1

@@ 2.5. Implications for model evaluation
The theoretical results above imply a sharp operational test: a model performs correct Bayesian inference if and only if its predictive entropy tracks (5) or (7) at every position.
@@ 3.3. Architectures
Parameter counts match transformers within 1%. These MLPs serve as controls testing whether hierarchical attention is architecturally necessary.
@@ 4.1. Bijection Wind Tunnel: Exact Hypothesis Elimination
Aggregate calibration could hide averaging artifacts. Figure 2 plots eight individual entropy trajectories. Each displays the characteristic staircase pattern: entropy drops discretely whenever a new input–output pair eliminates hypotheses, and collapses to near zero when an input repeats and the mapping is known. The model performs stepwise Bayesian elimination; it is not matching the curve in expectation but reproducing it sequence by sequence.
@@ 6.4. Architectural Necessity and MLP Failure
Transformers succeed because attention supplies the geometric mechanisms—-orthogonal bases, selective routing, and progressive focusing—-that MLPs lack. The failure of matched MLPs therefore serves as a clean demonstration that attention is not merely useful but architecturally necessary for Bayesian structure learning in context.
@@ 7.1. Bayesian Interpretations of Deep Learning
A long line of work interprets neural networks through a Bayesian lens, from classical analyses of predictive uncertainty (mackay1992bayesian; neal2012bayesian) to variational or stochastic approximations of posterior inference (graves2011practical; blundell2015weight). Recent papers argue that, in large-data limits, minimizing cross-entropy implicitly targets the Bayesian posterior predictive (xie2022explanation; von2023transformers). These results concern what training should produce at the population level. Our contribution is complementary: a controlled setting in which the true posterior is known, memorization is impossible, and one can directly test whether a finite transformer actually realizes this Bayesian computation.
@@ 7.2. In-Context Learning and Algorithmic Generalization
Transformers have been shown to perform algorithmic tasks in context, including arithmetic (garg2022can), synthetic induction (elman1990finding), and more general pattern extrapolation (akyurek2022learning; olsson2022context). Behaviorally, these models often resemble Bayesian learners, an observation formalized by recent explanatory theories (xie2022explanation; von2023transformers). However, prior work cannot distinguish true Bayesian computation from learned heuristics or memorized templates, because the ground-truth posterior is unknown for natural language tasks.
@@ 7.3. Mechanistic Interpretability and Attention Geometry
Mechanistic studies of transformers have revealed specialized attention heads for induction, copying, and retrieval (elhage2021mathematical; nanda2023progress). Other work has examined QKV spaces, circuit decomposition, and sparse structures that arise during training (olsson2022context). These studies provide qualitative and circuit-level insight into model behaviors.
@@ 7.4. Architectural Comparisons
Alternative sequence models—-state-space architectures (gu2022efficiently; gu2024mamba), convolutional variants (poli2023hyena), and deep MLPs—- often match transformers in perplexity on natural text. But perplexity conflates modeling and inference capability. Our results provide a finer test: whether an architecture can reproduce an analytic Bayesian posterior under strict non-memorization constraints. The capacity-matched MLP controls clarify that attention-based routing is not merely beneficial but necessary for Bayesian structure learning in context.
@@ 7.5. Training Dynamics
Finally, concurrent work analyzes the gradient dynamics that create these structures during training (dalal2024gradient). They show that attention and value updates follow coupled laws that produce a stable routing frame and a progressively refined value manifold. Our empirical findings align with this picture: attention stabilizes early, while value vectors continue to encode the posterior with increasing resolution. Together, these perspectives connect the optimization trajectory to the geometric structure that implements Bayesian inference.
@@ 9. Conclusion
Transformers succeed here not because they are large, but because their architecture furnishes the right inductive bias: residual streams that carry evolving belief states, attention that routes information selectively, and feed-forward layers that implement local Bayesian updates. Together, these components carve a Bayesian manifold inside the model’s representation space. Understanding how this manifold emerges, scales, and ultimately degrades in real-world language remains an important direction for future work.

### v2

@@ 1. Introduction
Clarification on “Bayesian inference.” We do not claim a Bayesian posterior over network weights; we show that the learned predictor implements the Bayesian posterior predictive over latent task variables—the filtering posterior over hidden states in HMMs, or the elimination posterior over bijections. This is a statement about the input-output function the transformer computes, not about weight-space uncertainty.
@@ 2.5. Implications for model evaluation
The theoretical results above imply a practical diagnostic: a model that achieves the correct posterior entropy at every position is functionally Bayesian—it produces predictions with the same uncertainty profile as the exact posterior. [^1] Entropy-matching is necessary but not sufficient for distributional convergence in general, since different distributions can share the same entropy. However, direct verification via KL divergence confirms that our trained transformers match the full Bayesian posterior: $\text{KL}(\text{model}\|\text{Bayes})\approx 0$ nats and total variation distance $3\%$ for new keys, with exact retrieval (KL =0=0, TVD =0=0) for repeated keys in the with-replacement setting. Combined with the cross-entropy training objective (whose unique population minimizer is the Bayesian posterior predictive), low entropy calibration error provides strong evidence for Bayesian computation.
@@ 3.3. Architectures
Parameter counts match transformers within 1%. These MLPs serve as controls testing whether hierarchical attention is essential for the task.
@@ 4.1. Bijection Wind Tunnel: Exact Hypothesis Elimination
Aggregate calibration could hide averaging artifacts. Figure 2 plots eight individual entropy trajectories. Each displays the characteristic staircase pattern: entropy drops discretely whenever a new input–output pair eliminates hypotheses, and collapses to near zero when an input repeats and the mapping is known. The model performs stepwise Bayesian elimination, reproducing the curve sequence by sequence rather than merely matching it in expectation.
@@ 6.4. Architectural Necessity and MLP Failure
Transformers succeed because attention supplies the geometric mechanisms—-orthogonal bases, selective routing, and progressive focusing—-that MLPs lack. The failure of matched MLPs therefore serves as a clean demonstration that attention is essential for Bayesian structure learning in context, at least among the architectures tested.
@@ 7.1. Bayesian Interpretations of Deep Learning
A long line of work interprets neural networks through a Bayesian lens, from classical analyses of predictive uncertainty (MacKay, 1992; Neal, 2012) to variational or stochastic approximations of posterior inference (Graves, 2011; Blundell et al., 2015). Recent papers argue that, in large-data limits, minimizing cross-entropy implicitly targets the Bayesian posterior predictive (Xie et al., 2022; von Oswald et al., 2023). These results concern what training should produce at the population level. Our contribution is complementary: a controlled setting in which the true posterior is known, memorization is impossible, and one can directly test whether a finite transformer actually realizes this Bayesian computation.
@@ 7.2. In-Context Learning and Algorithmic Generalization
Transformers have been shown to perform algorithmic tasks in context, including arithmetic (Garg et al., 2022), synthetic induction (Elman, 1990), and more general pattern extrapolation (Akyürek and Andreas, 2022; Olsson et al., 2022). Behaviorally, these models often resemble Bayesian learners, an observation formalized by recent explanatory theories (Xie et al., 2022; von Oswald et al., 2023). However, prior work cannot distinguish true Bayesian computation from learned heuristics or memorized templates, because the ground-truth posterior is unknown for natural language tasks.
@@ 7.3. Mechanistic Interpretability and Attention Geometry
Mechanistic studies of transformers have revealed specialized attention heads for induction, copying, and retrieval (Elhage et al., 2021; Nanda et al., 2023). Other work has examined QKV spaces, circuit decomposition, and sparse structures that arise during training (Olsson et al., 2022). These studies provide qualitative and circuit-level insight into model behaviors.
@@ 7.4. Architectural Comparisons
Alternative sequence models—-state-space architectures (Gu et al., 2022; Gu and Dao, 2023), convolutional variants (Poli et al., 2023), and deep MLPs—- often match transformers in perplexity on natural text. But perplexity conflates modeling and inference capability. Our results provide a finer test: whether an architecture can reproduce an analytic Bayesian posterior under strict non-memorization constraints. The capacity-matched MLP controls clarify that attention-based routing is essential for Bayesian structure learning in context, at least relative to flat feed-forward architectures.
@@ 7.5. Training Dynamics
Finally, concurrent work analyzes the gradient dynamics that create these structures during training (Aggarwal et al., 2025). They show that attention and value updates follow coupled laws that produce a stable routing frame and a progressively refined value manifold. Our empirical findings align with this picture: attention stabilizes early, while value vectors continue to encode the posterior with increasing resolution. Together, these perspectives connect the optimization trajectory to the geometric structure that implements Bayesian inference.
@@ 9. Conclusion
Transformers succeed here because their architecture furnishes the right inductive bias—not because of scale: residual streams that carry evolving belief states, attention that routes information selectively, and feed-forward layers that implement local Bayesian updates. Together, these components carve a Bayesian manifold inside the model’s representation space. Understanding how this manifold emerges, scales, and ultimately degrades in real-world language remains an important direction for future work.